OpenStudy (anonymous):

Aliyah1981 0 If a line on a graph is approximately the distance (y) in feet a person walks in (x) hours. What does the slope of the line represent? And how is the graph usefull? a few moments ago

OpenStudy (e.mccormick):

First, have you heard a few different definitions of slope?

OpenStudy (anonymous):

No

OpenStudy (e.mccormick):

Ever seen:\[\frac{rise}{run}\;\mathrm{or}\;\frac{\Delta y}{\Delta x}\]

OpenStudy (anonymous):

Yes

OpenStudy (e.mccormick):

Here, the second one is important, \(\frac{\Delta y}{\Delta x}\). Do you know what the delta \((\Delta)\) part of that means?

OpenStudy (anonymous):

No

OpenStudy (e.mccormick):

Ah, they chose Delta for Difference, the Difference in Y over the Difference in X. However, this is more commonly called the Change in Y over the Change in X. So, if you think about it as the changes, what are Y and X in your problem. Put them in there verbally as "The change in...."

OpenStudy (e.mccormick):

Here is a related concep: Ever heard of a car being rated in MPH?

OpenStudy (anonymous):

No

OpenStudy (e.mccormick):

Miles Per Hour, where Per is another way to say Divided By.

OpenStudy (anonymous):

Ok

OpenStudy (e.mccormick):

That is the speed of a car, MPH. So, with that in mind, take another look at your problem.

OpenStudy (e.mccormick):

Any ideas yet?

OpenStudy (anonymous):

No

OpenStudy (e.mccormick):

Well, there is no car in this... just a person.... and there is no driving, just walkng... so, if I say they are somehow similar, there is some sort of relationship there.